OpenStudy (anonymous):

simplify the expression 5y/y(2y-1)+2/2y-1

OpenStudy (anonymous):

\[\frac{ 5y }{ y(2y-1) } + \frac{2}{2y-1}\] You can simplify the right side to \[\frac{ 5 }{ 2y-1 } + \frac{2}{2y-1}\] Add the two fractions \[\frac{ 7 }{ 2y-1 }\]